OpenStudy (anonymous):

This is due tommorow cuz i wasnt at school yesterday & there was a sub & i didnt get a chance to be taught.PLEASE..... 1) y-7+4y=13 y= 2) 3(a-5)+19=-2 a= 3) -9-3(2r-1)=-18 r= 4) (x-3)-(x-5)+ 2x=0 x= 5) 5(g-7)=2[g-5]=0 g= Please Helpp:)

OpenStudy (anonymous):

PLease show ur work i really will appreciate it:D

OpenStudy (anonymous):

thx for the medal:)

OpenStudy (anonymous):

haha no problem, you didnt have any :D

OpenStudy (anonymous):

oh i knew that lol:)

OpenStudy (anonymous):

haha i feel lonly, can i get onee?

OpenStudy (anonymous):

1)y-7+4y=13 implies 5y-7=13 implies 5y-7+7=13+7 implies 5y=20 implies 5y/5=20/5 implies y=4

OpenStudy (anonymous):

ill be rit back gootta go do sumthing:)